# Chapter 4 Discrete subgroups of semisimple Lie groups, beyond lattices

### 26 Feb 2024 v1

@@ Abstract
Chapter 4 Discrete subgroups of semisimple Lie groups, beyond lattices
@@ 1 Introduction
We say that a Lie group $G$ is simple if its Lie algebra is simple, i.e. nonabelian with no nonzero proper ideals; equivalently, all infinite closed normal subgroups of $G$ have finite index in $G$ and are nonabelian. Simple Lie algebras have been completely classified by É. Cartan, leading to a classification of simple Lie groups up to local isomorphism. (Recall that two Lie groups $G_{1}$ and $G_{2}$ are said to be locally isomorphic if they have the same Lie algebra; equivalently, some finite cover of the identity component of $G_{1}$ is isomorphic to some finite cover of the identity component of $G_{2}$.) Noncompact simple Lie groups come in several infinite families, given in Table 1, and 17 (up to local isomorphism) additional groups, called exceptional (see e.g. [88, Ch. X]).
@@ D III
We say that a Lie group $G$ is semisimple if it is locally isomorphic to a direct product $G_{1}\times\dots\times G_{\ell}$ of simple Lie groups $G_{i}$, called the simple factors of $G$; in that case, if $G$ is connected and simply connected, then it is actually isomorphic to such a direct product $G_{1}\times\dots\times G_{\ell}$. For instance, $\mathrm{SO}(2,2)$ and $\mathrm{SO}(4,\mathbb{C})$ are semisimple (they are locally isomorphic to $\mathrm{PSL}(2,\mathbb{R})\times\mathrm{PSL}(2,\mathbb{R})$ and $\mathrm{PSL}(2,\mathbb{C})\times\mathrm{PSL}(2,\mathbb{C})$, respectively). Any connected semisimple Lie group is the identity component (for the real topology) of the real points of some $\mathbb{R}$-algebraic group (see [44, § 2.14]).
@@
These properties are typically invariant under replacing $\Gamma$ by a finite-index subgroup. This will allow us to sometimes reduce to torsion-free $\Gamma$: indeed, the Selberg lemma [123, Lem. 8] states that any finitely generated subgroup of $G$ admits a finite-index subgroup which is torsion-free.
@@ 2.2 Examples
and the Lie group $G:=\mathrm{SO}(J_{p,q},\mathbb{R})\simeq\mathrm{SO}(p,q)$. Then $\Gamma:=G\cap\mathrm{SL}(n,\mathbb{Z}[\sqrt{2}])$ is a cocompact lattice in $G$.
@@
In both Examples 4.2 and 4.3, the group $\Gamma$ is arithmetic in $G$, i.e. there is a homomorphism $\pi:\mathbf{H}\to\mathbf{G}$ of semisimple $\mathbb{Q}$-algebraic groups such that $G=\mathbf{G}_{\mathbb{R}}$, such that the kernel of $\pi$ in $\mathbf{H}_{\mathbb{R}}$ is compact, and such that $\Gamma$ is commensurable to $\pi(\mathbf{H}_{\mathbb{Z}})$ (see [133]).
@@
On the other hand, in noncompact simple Lie groups which are not locally isomorphic to $\mathrm{SO}(n,1)$ or $\mathrm{SU}(n,1)$, all lattices are arithmetic (as a consequence of superrigidity, see Section 2.4).
2.3 Rank one versus higher rank
@@
The real rank of $G$, denoted $\mathrm{rank}_{\mathbb{R}}(G)$, is the maximum dimension of a subgroup of $G$ which is diagonalisable over $\mathbb{R}$; equivalently, for noncompact $G$, it is the maximum dimension of a totally geodesic subspace of the Riemannian symmetric space $G/K$ which is flat (i.e. of constant zero sectional curvature).
@@
  • any noncocompact lattice $\Gamma$ of $G$ is relatively hyperbolic with respect to some collection $\mathcal{P}$ of subgroups which are virtually (i.e. up to finite index) nilpotent: this means that $\Gamma$ acts properly discontinuously by isometries on some visual Gromov hyperbolic proper metric space $X$ (e.g. $X=G/K$), and with compact quotient on some closed subset of $X$ of the form $X\smallsetminus\bigcup_{P\in\mathcal{P}}B_{P}$ where each $B_{P}$ is a $P$-invariant open horoball of $X$ and $B_{P}\cap B_{P^{\prime}}=\emptyset$ for $P\neq P^{\prime}$ (see [87, § 4] for details).
@@
More generally, lattices $\Gamma$ in simple Lie groups $G$ with $\mathrm{rank}_{\mathbb{R}}(G)\geq\nolinebreak 2$ tend to have global fixed points when they act on various classes of spaces. For instance, any continuous action by affine isometries of $\Gamma$ on a Hilbert space has a global fixed point. This property, which is equivalent to Kazhdan’s property (T) (see [133, Ch. 13]), is also satisfied by lattices in the rank-one Lie groups $\mathrm{Sp}(n,1)$ with $n\geq 2$ or $F_{4(-20)}$. However, other fixed point properties actually distinguish higher rank from rank one. For instance, for any simple Lie group $G$ with $\mathrm{rank}_{\mathbb{R}}(G)\geq 2$ and any $\sigma$-finite positive measure $\nu$ on a standard Borel space, any continuous affine action of a lattice of $G$ on $L^{p}(\nu)$ for $1<p<+\infty$ has a global fixed point, by Bader–Furman–Gelander–Monod; on the other hand, by Pansu and Bourdon–Pajot, any cocompact lattice $\Gamma$ in a simple Lie group $G$ with $\mathrm{rank}_{\mathbb{R}}(G)=\nolinebreak 1$ (and more generally, any Gromov hyperbolic group $\Gamma$) admits fixed-point-free affine isometric actions on $L^{p}(\Gamma)$ whose linear part is the regular representation, for any $p>1$ large enough. See [4].
@@
Note that for an irreducible lattice $\Gamma$ of $G$, the finite normal subgroups of $\Gamma$ are easily described as the subgroups of the finite abelian group $\Gamma\cap Z(G)$ (using Borel’s Density Theorem [28]). On the other hand, much more effort is required to understand the finite-index normal subgroups of $\Gamma$. By [11], for $\Gamma=\mathrm{SL}(n,\mathbb{Z})$ with $n\geq 3$, any finite-index normal subgroup of $\Gamma$ is a congruence subgroup, i.e. contains the kernel of the natural projection $\mathrm{SL}(n,\mathbb{Z})\to\mathrm{SL}(n,\mathbb{Z}/m\mathbb{Z})$ for some $m\geq 1$; this is false for $\Gamma=\mathrm{SL}(2,\mathbb{Z})$. In general, it is conjectured that lattices of $G$ have a slightly weaker form of this “Congruence Subgroup Property” if and only if $\mathrm{rank}_{\mathbb{R}}(G)\geq 2$: see [125].
@@ 2.4 Deformations and rigidity
By a continuous deformation of $\Gamma$ in $G$ we mean a continuous path $(\rho_{t})_{t\in[0,1)}$ in $\mathrm{Hom}(\Gamma,G)$ where $\rho_{0}$ is the natural inclusion of $\Gamma$ in $G$. Certain continuous deformations of $\Gamma$ in $G$ are considered trivial: namely, those of the form $\rho_{t}=g_{t}\,\rho_{0}(\cdot)\,g_{t}^{-1}$ where $(g_{t})_{t\in[0,1)}$ is a continuous path in $G$ (and $g_{0}$ is the identity element). In other words, if $\mathrm{Hom}(\Gamma,G)/G$ denotes the quotient of $\mathrm{Hom}(\Gamma,G)$ by the natural action of $G$ by conjugation at the target, then the trivial deformations are those whose image in $\mathrm{Hom}(\Gamma,G)/G$ is constant.
@@
This implies (see Section 2.1) that the fundamental group of any locally symmetric space $\Gamma\backslash G/K$ completely determines its geometry.
Margulis superrigidity (Margulis, Corlette, Gromov–Schoen, see e.g. [133, Th. 16.1.4]).
@@
Here “$\rho$ continuously extends to $G$ up to finite index and to bounded error” means that there exist a finite-index subgroup $\Gamma^{\prime}$ of $\Gamma$, a continuous homomorphism $\rho_{G}:G\to\mathrm{GL}(d,\mathbb{R})$, and a compact subgroup $C$ of $\mathrm{GL}(d,\mathbb{R})$ centralising $\rho_{G}(G)$ such that $\rho(\gamma)\in\rho_{G}(\gamma)C$ for all $\gamma\in\Gamma^{\prime}$. Under an appropriate assumption on the image of $\rho$, we can take $C$ to be trivial. “Algebraically simply connected” is a technical assumption which is always satisfied up to passing to a finite cover: see [133, § 16.1].
Margulis used his superrigidity (over $\mathbb{R}$ as above, but also over non-Archimedean local fields) to prove that if $G$ is semisimple with no compact simple factors and if $\mathrm{rank}_{\mathbb{R}}(G)\geq 2$, then all irreducible lattices $\Gamma$ of $G$ are arithmetic in the sense of Section 2.2. The same conclusion holds when $G$ is locally isomorphic to $\mathrm{Sp}(n,1)$ with $n\geq 2$ or $F_{4(-20)}$, as superrigidity holds for these rank-one groups as well.
@@ 3 A change of paradigm
To be more precise, we are interested in infinite discrete subgroups $\Gamma$ of semisimple Lie groups $G$ that admit continuous deformations $(\rho_{t})_{t\in[0,1)}\linebreak\subset\mathrm{Hom}(\Gamma,G)$ as in Section 2.4 which, not only are nontrivial, but also satisfy that each $\rho_{t}$ is injective with discrete image, so that the $\rho_{t}(\Gamma)$ for $t>0$ are still discrete subgroups of $G$ isomorphic (but not conjugate) to $\Gamma$. An ideal situation is when the natural inclusion $\rho_{0}:\Gamma\hookrightarrow G$ admits a full open neighbourhood in $\mathrm{Hom}(\Gamma,G)$ consisting entirely of injective and discrete representations, with a nonconstant image in $\mathrm{Hom}(\Gamma,G)/G$.
We are thus led, for given discrete subgroups $\Gamma$ of $G$, to study subsets of $\mathrm{Hom}(\Gamma,G)$ consisting of injective and discrete representations, and their images in the corresponding character varieties. In this framework, we discuss so-called higher Teichmüller theory in Section 3.4 below.
@@
3.1 Examples in real rank one
@@ 3.1.1 Schottky groups
Let $\mathcal{D}:=\partial_{\infty}X\smallsetminus\bigcup_{i=1}^{m}(B_{i}^{-}\cup B_{i}^{+})$ and $\Omega:=\mathrm{Int}(\bigcup_{\gamma\in\Gamma}\gamma\cdot\nolinebreak\mathcal{D})$. Then $\Omega$ is an open subset of $\partial_{\infty}X$ on which $\Gamma$ acts properly discontinuously with fundamental domain $\mathcal{D}$.
@@
Since Schottky groups $\Gamma$ are nonabelian free groups, they admit, as in Section 2.4, many nontrivial continuous deformations $(\rho_{t})_{t\in[0,1)}\subset\mathrm{Hom}(\Gamma,G)$, obtained by independently deforming the image of each generator $\gamma_{i}$. Some of these deformations $(\rho_{t})_{t\in[0,1)}$ are “good” in the sense that for every $t\in[0,1)$, the group $\rho_{t}(\Gamma)$ still has a ping pong configuration analogous to that of $\Gamma$, hence $\rho_{t}$ is injective with discrete image by arguing as in Claim 4.6. If the open disks $B_{1}^{\pm},\dots,B_{m}^{\pm}$ in the initial configuration have pairwise disjoint closures (i.e. $\Gamma$ is a “strong” Schottky group), then all small deformations are “good”: the natural inclusion $\rho_{0}:\Gamma\hookrightarrow G$ admits an open neighbourhood in $\mathrm{Hom}(\Gamma,G)$ consisting entirely of injective and discrete representations, with a ping pong configuration analogous to that of $\Gamma$.
3.1.2 Quasi-Fuchsian groups
@@
Let $S$ be a closed orientable surface of genus $g\geq 2$. By the Uniformisation Theorem (see Section 1), there exist injective and discrete representations from the fundamental group $\pi_{1}(S)$ to $\mathrm{PSL}(2,\mathbb{R})$. These representations form two connected components of $\mathrm{Hom}(\pi_{1}(S),\mathrm{PSL}(2,\mathbb{R}))$, switched by conjugation by elements of $\mathrm{PGL}(2,\mathbb{R})\smallsetminus\mathrm{PSL}(2,\mathbb{R})$ (i.e. by orientation-reversing isometries of $\mathbb{H}^{2}$). The image of either of these connected components in $\mathrm{Hom}(\pi_{1}(S),\mathrm{PSL}(2,\mathbb{R}))/\mathrm{PSL}(2,\mathbb{R})$ identifies with the Teichmüller space of $S$, which is homeomorphic to $\mathbb{R}^{6g-6}$.
@@
3.1.3 Deformations of Fuchsian representations for higher-dimensional groups
Recall that $\mathrm{PSL}(2,\mathbb{R})\simeq\mathrm{SO}(2,1)_{0}$ and $\mathrm{PSL}(2,\mathbb{C})\simeq\mathrm{SO}(3,1)_{0}$. We now consider any integer $n\geq 2$ and let $\Gamma$ be a cocompact lattice of $\mathrm{SO}(n,1)_{0}$. As above, we can see $\Gamma$ as a discrete subgroup of $\mathrm{SO}(n+1,1)$ (which is not a lattice anymore). Interestingly, although all continuous deformations of $\Gamma$ in $\mathrm{SO}(n,1)$ are trivial for $n\geq 3$ (by Mostow rigidity, see Section 2.4), there can exist nontrivial continuous deformations of $\Gamma$ in $\mathrm{SO}(n+1,1)$. Such deformations were constructed in [92, 106] based on a construction of Thurston called bending.
@@ 3.1.3 Deformations of Fuchsian representations for higher-dimensional groups
In either case, Johnson and Millson [92] observed that for small enough $t>0$ the representation $\rho_{t}$ has Zariski-dense image in $\mathrm{SO}(n+1,1)$; moreover, $\rho_{t}$ is still injective and discrete for small $t$ (see Section 4.2).
@@
    The fact that $\rho_{t}$ is injective and discrete for small $t$ also follows from Maskit’s combination theorems, which generalise the idea of ping pong to amalgamated free products and HNN extensions (see [111, § VIII.E.3]).
3.2 Ping pong in higher real rank
Examples of “flexible” discrete subgroups of higher-rank semisimple Lie groups $G$ which are nonabelian free groups can be constructed by generalising the classical Schottky groups of Section 3.1 in various ways. Let us mention three geometric constructions.
3.2.1 Ping pong in projective space

### 28 Mar 2024 v2

# Chapter 0 Discrete subgroups of semisimple Lie groups, beyond lattices

@@ Abstract
Chapter 0 Discrete subgroups of semisimple Lie groups, beyond lattices
@@ 1 Introduction
We say that a Lie group $G$ is simple if its Lie algebra is simple, i.e. nonabelian with no nonzero proper ideals; equivalently, all infinite closed normal subgroups of $G$ have finite index in $G$ and are nonabelian. Simple Lie algebras have been completely classified by É. Cartan, leading to a classification of simple Lie groups up to local isomorphism. (Recall that two Lie groups $G_{1}$ and $G_{2}$ are said to be locally isomorphic if they have the same Lie algebra; equivalently, some finite cover of the identity component of $G_{1}$ is isomorphic to some finite cover of the identity component of $G_{2}$.) Noncompact simple Lie groups come in several infinite families, given in Table 1, and 17 (up to local isomorphism) additional groups, called exceptional (see e.g. [90, Ch. X]).
@@ D III
We say that a Lie group $G$ is semisimple if it is locally isomorphic to a direct product $G_{1}\times\dots\times G_{\ell}$ of simple Lie groups $G_{i}$, called the simple factors of $G$; in that case, if $G$ is connected and simply connected, then it is actually isomorphic to such a direct product $G_{1}\times\dots\times G_{\ell}$. For instance, $\mathrm{SO}(2,2)$ and $\mathrm{SO}(4,\mathbb{C})$ are semisimple (they are locally isomorphic to $\mathrm{PSL}(2,\mathbb{R})\times\mathrm{PSL}(2,\mathbb{R})$ and $\mathrm{PSL}(2,\mathbb{C})\times\mathrm{PSL}(2,\mathbb{C})$, respectively). Any connected semisimple Lie group is the identity component (for the real topology) of the real points of some $\mathbb{R}$-algebraic group (see [45, § 2.14]).
@@
These properties are typically invariant under replacing $\Gamma$ by a finite-index subgroup. This will allow us to sometimes reduce to torsion-free $\Gamma$: indeed, the Selberg lemma [125, Lem. 8] states that any finitely generated subgroup of $G$ admits a finite-index subgroup which is torsion-free.
@@ 2 Examples
and the Lie group $G:=\mathrm{SO}(J_{p,q},\mathbb{R})\simeq\mathrm{SO}(p,q)$. Then $\Gamma:=\linebreak G\cap\mathrm{SL}(n,\mathbb{Z}[\sqrt{2}])$ is a cocompact lattice in $G$.
@@
In both Examples 2.2 and 2.3, the group $\Gamma$ is arithmetic in $G$, i.e. there is a homomorphism $\pi:\mathbf{H}\to\mathbf{G}$ of semisimple $\mathbb{Q}$-algebraic groups such that $G=\mathbf{G}_{\mathbb{R}}$, such that the kernel of $\pi$ in $\mathbf{H}_{\mathbb{R}}$ is compact, and such that $\Gamma$ is commensurable to $\pi(\mathbf{H}_{\mathbb{Z}})$ (see [135]).
@@
On the other hand, in noncompact simple Lie groups which are not locally isomorphic to $\mathrm{SO}(n,1)$ or $\mathrm{SU}(n,1)$, all lattices are arithmetic (as a consequence of superrigidity, see Section 4).
3 Rank one versus higher rank
@@
The real rank of $G$, denoted $\mathrm{rank}_{\mathbb{R}}(G)$, is the maximum dimension of a closed connected subgroup of $G$ which is diagonalisable over $\mathbb{R}$; equivalently, for noncompact $G$, it is the maximum dimension of a totally geodesic subspace of the Riemannian symmetric space $G/K$ which is flat (i.e. of constant zero sectional curvature).
@@
  • any noncocompact lattice $\Gamma$ of $G$ is relatively hyperbolic with respect to some collection $\mathcal{P}$ of subgroups which are virtually (i.e. up to finite index) nilpotent: this means that $\Gamma$ acts properly discontinuously by isometries on some visual Gromov hyperbolic proper metric space $X$ (e.g. $X=G/K$), and with compact quotient on some closed subset of $X$ of the form $X\smallsetminus\bigcup_{P\in\mathcal{P}}B_{P}$ where each $B_{P}$ is a $P$-invariant open horoball of $X$ and $B_{P}\cap B_{P^{\prime}}=\emptyset$ for $P\neq P^{\prime}$ (see [89, § 4] for details).
@@
More generally, lattices $\Gamma$ in simple Lie groups $G$ with $\mathrm{rank}_{\mathbb{R}}(G)\geq\nolinebreak 2$ tend to have global fixed points when they act on various classes of spaces. For instance, any continuous action by affine isometries of $\Gamma$ on a Hilbert space has a global fixed point. This property, which is equivalent to Kazhdan’s property (T) (see [135, Ch. 13]), is also satisfied by lattices in the rank-one Lie groups $\mathrm{Sp}(n,1)$ with $n\geq 2$ or $F_{4(-20)}$. However, other fixed point properties actually distinguish higher rank from rank one. For instance, for any simple Lie group $G$ with $\mathrm{rank}_{\mathbb{R}}(G)\geq 2$ and any $\sigma$-finite positive measure $\nu$ on a standard Borel space, any continuous affine isometric action of a lattice of $G$ on $L^{p}(\nu)$ for $1<p<+\infty$ has a global fixed point, by Bader–Furman–Gelander–Monod; on the other hand, by Pansu and Bourdon–Pajot, any cocompact lattice $\Gamma$ in a simple Lie group $G$ with $\mathrm{rank}_{\mathbb{R}}(G)=\nolinebreak 1$ (and more generally, any Gromov hyperbolic group $\Gamma$) admits fixed-point-free affine isometric actions on $L^{p}(\Gamma)$ whose linear part is the regular representation, for any $p>1$ large enough. See [4].
@@
Note that for an irreducible lattice $\Gamma$ of $G$, the finite normal subgroups of $\Gamma$ are easy to describe: for connected $G$, they are the subgroups of the finite abelian group $\Gamma\cap Z(G)$ (using Borel’s Density Theorem [28]). On the other hand, much more effort is required to understand the finite-index normal subgroups of $\Gamma$. By [11], for $\Gamma=\mathrm{SL}(n,\mathbb{Z})$ with $n\geq 3$, any finite-index normal subgroup of $\Gamma$ is a congruence subgroup, i.e. contains the kernel of the natural projection $\mathrm{SL}(n,\mathbb{Z})\to\mathrm{SL}(n,\mathbb{Z}/m\mathbb{Z})$ for some $m\geq 1$; this is false for $\Gamma=\mathrm{SL}(2,\mathbb{Z})$. In general, it is conjectured that lattices of $G$ have a slightly weaker form of this “Congruence Subgroup Property” if and only if $\mathrm{rank}_{\mathbb{R}}(G)\geq 2$: see [127].
@@ 4 Deformations and rigidity
By a continuous deformation of $\Gamma$ in $G$ we mean a continuous path $(\rho_{t})_{t\in[0,1)}$ in $\mathrm{Hom}(\Gamma,G)$ where $\rho_{0}$ is the natural inclusion of $\Gamma$ in $G$. Certain continuous deformations of $\Gamma$ in $G$ are considered trivial: namely, those of the form $\rho_{t}=g_{t}\,\rho_{0}(\cdot)\,g_{t}^{-1}$ where $(g_{t})_{t\in[0,1)}$ is a continuous path in $G$ (and $g_{0}$ is the identity element). In other words, if $\mathrm{Hom}(\Gamma,G)/G$ denotes the quotient of $\mathrm{Hom}(\Gamma,G)$ by the natural action of $G$ by conjugation at the target, then the trivial deformations are those whose image in $\mathrm{Hom}(\Gamma,G)/G$ are constant.
@@
This implies (see Section 1) that the fundamental group of any locally symmetric space $\Gamma\backslash G/K$ completely determines its geometry.
Margulis superrigidity (Margulis, Corlette, Gromov–Schoen, see e.g. [135, Th. 16.1.4]).
@@
Here “$\rho$ continuously extends to $G$ up to finite index and to bounded error” means that there exist a finite-index subgroup $\Gamma^{\prime}$ of $\Gamma$, a continuous homomorphism $\rho_{G}:G\to\mathrm{GL}(d,\mathbb{R})$, and a compact subgroup $C$ of $\mathrm{GL}(d,\mathbb{R})$ centralising $\rho_{G}(G)$ such that $\rho(\gamma)\in\rho_{G}(\gamma)C$ for all $\gamma\in\Gamma^{\prime}$. Under an appropriate assumption on the image of $\rho$, we can take $C$ to be trivial. “Algebraically simply connected” is a technical assumption which is always satisfied up to passing to a finite cover: see [135, § 16.1].
Margulis used his superrigidity (over $\mathbb{R}$ as above, but also over non-Archimedean local fields) to prove that if $G$ is semisimple with no compact simple factors and if $\mathrm{rank}_{\mathbb{R}}(G)\geq 2$, then all irreducible lattices $\Gamma$ of $G$ are arithmetic in the sense of Section 2. The same conclusion holds when $G$ is locally isomorphic to $\mathrm{Sp}(n,1)$ with $n\geq 2$ or $F_{4(-20)}$, as superrigidity holds for these rank-one groups as well.
@@ 3 A change of paradigm
To be more precise, we are interested in infinite discrete subgroups $\Gamma$ of semisimple Lie groups $G$ that admit continuous deformations $(\rho_{t})_{t\in[0,1)}\linebreak\subset\mathrm{Hom}(\Gamma,G)$ as in Section 4 which, not only are nontrivial, but also satisfy that each $\rho_{t}$ is injective with discrete image, so that the $\rho_{t}(\Gamma)$ for $t>0$ are still discrete subgroups of $G$ isomorphic (but not conjugate) to $\Gamma$. An ideal situation is when the natural inclusion $\rho_{0}:\Gamma\hookrightarrow G$ admits a full open neighbourhood in $\mathrm{Hom}(\Gamma,G)$ consisting entirely of injective and discrete representations, with a nonconstant image in $\mathrm{Hom}(\Gamma,G)/G$.
We are thus led, for given discrete subgroups $\Gamma$ of $G$, to study subsets of $\mathrm{Hom}(\Gamma,G)$ consisting of injective and discrete representations, and their images in the corresponding character varieties. In this framework, we discuss so-called higher Teichmüller theory in Section 4 below.
@@
1 Examples in real rank one
@@
Let $\mathcal{D}:=\partial_{\infty}X\smallsetminus\bigcup_{i=1}^{m}(B_{i}^{-}\cup B_{i}^{+})$ and $\Omega:=\linebreak\mathrm{Int}(\bigcup_{\gamma\in\Gamma}\gamma\cdot\nolinebreak\mathcal{D})$. Then $\Omega$ is an open subset of $\partial_{\infty}X$ on which $\Gamma$ acts properly discontinuously with fundamental domain $\mathcal{D}$.
@@
Since Schottky groups $\Gamma$ are nonabelian free groups, they admit, as in Section 4, many nontrivial continuous deformations $(\rho_{t})_{t\in[0,1)}\subset\mathrm{Hom}(\Gamma,G)$, obtained by independently deforming the image of each generator $\gamma_{i}$. Some of these deformations $(\rho_{t})_{t\in[0,1)}$ are “good” in the sense that for every $t\in[0,1)$, the group $\rho_{t}(\Gamma)$ still has a ping pong configuration analogous to that of $\Gamma$, hence $\rho_{t}$ is injective with discrete image by arguing as in Claim 3.2. If the open disks $B_{1}^{\pm},\dots,B_{m}^{\pm}$ in the initial configuration have pairwise disjoint closures (i.e. $\Gamma$ is a “strong” Schottky group), then all small deformations are “good”: the natural inclusion $\rho_{0}:\Gamma\hookrightarrow G$ admits an open neighbourhood in $\mathrm{Hom}(\Gamma,G)$ consisting entirely of injective and discrete representations, with a ping pong configuration analogous to that of $\Gamma$.
Quasi-Fuchsian groups
@@
Let $S$ be a closed orientable surface of genus $g\geq 2$. By the Uniformisation Theorem (see Section 1), there exist injective and discrete representations from the fundamental group $\pi_{1}(S)$ to $\mathrm{PSL}(2,\mathbb{R})$. These representations form two connected components of $\mathrm{Hom}(\pi_{1}(S),\mathrm{PSL}(2,\mathbb{R}))$ [77], switched by conjugation by elements of $\mathrm{PGL}(2,\mathbb{R})\smallsetminus\mathrm{PSL}(2,\mathbb{R})$ (i.e. by orientation-reversing isometries of $\mathbb{H}^{2}$). The image of either of these connected components in $\mathrm{Hom}(\pi_{1}(S),\mathrm{PSL}(2,\mathbb{R}))/\mathrm{PSL}(2,\mathbb{R})$ identifies with the Teichmüller space of $S$, which is homeomorphic to $\mathbb{R}^{6g-6}$.
@@
Deformations of Fuchsian representations for higher-dimensional groups
Recall that $\mathrm{PSL}(2,\mathbb{R})\simeq\mathrm{SO}(2,1)_{0}$ and $\mathrm{PSL}(2,\mathbb{C})\simeq\mathrm{SO}(3,1)_{0}$. We now consider any integer $n\geq 2$ and let $\Gamma$ be a cocompact lattice of $\mathrm{SO}(n,1)_{0}$. As above, we can see $\Gamma$ as a discrete subgroup of $\mathrm{SO}(n+1,1)$ (which is not a lattice anymore). Interestingly, although all continuous deformations of $\Gamma$ in $\mathrm{SO}(n,1)$ are trivial for $n\geq 3$ (by Mostow rigidity, see Section 4), there can exist nontrivial continuous deformations of $\Gamma$ in $\mathrm{SO}(n+1,1)$. Such deformations were constructed in [94, 108] based on a construction of Thurston called bending.
@@
In either case, Johnson and Millson [94] observed that for small enough $t>0$ the representation $\rho_{t}$ has Zariski-dense image in $\mathrm{SO}(n+1,1)$; moreover, $\rho_{t}$ is still injective and discrete for small $t$ (see Section 2).
@@
    The fact that $\rho_{t}$ is injective and discrete for small $t$ also follows from Maskit’s combination theorems, which generalise the idea of ping pong to amalgamated free products and HNN extensions (see [113, § VIII.E.3]).
2 Ping pong in higher real rank
Examples of “flexible” discrete subgroups of higher-rank semisimple Lie groups $G$ which are nonabelian free groups can be constructed by generalising the classical Schottky groups of Section 1 in various ways. Let us mention three geometric constructions.
Ping pong in projective space